OpenStudy (islinger12):

List the sides of ABC in order from shortest to longest, when A = 10x – 7, B = 7x – 10, and C = 47 – 2x.

OpenStudy (bironic):

add all the values together, so you'd get 15x +30. Then you minus 30, which makes 15x=-30. Divide 15 on both sides so you get x=-2. Then, plug in the x value for all sides. So A=10(-2) - 7, B=7(-2) - 10, and C=47 - 2(-2). If you do the math, you'll get A=-27, B=-24, and C=51. The answer is CBA.

OpenStudy (mathstudent55):

The solution above written by @bironic makes no sense. If A, B, and C are the angles of triangle ABC, then how can you have angles with negative measures? What you need to do is add the expressions for the measures of the angles and set the sum equal to 180. Then find x. Once you have x, use x in each expression and find the measure of each angle. The length of the sides is in the same order of size as the measure of the opposite angles. The longest side is opposite the largest angle. the next longest side is opposite the next largest angle.The smallest side is opposite the smallest angle.